OpenStudy (anonymous):

[3.05] Given the function g(x) = 4x − 5, find g(−3). −17 7 −7 2 I don't know how to solve this? Please explain

OpenStudy (anonymous):

The lesson on this is confusing

OpenStudy (anonymous):

g(x) = 4x − 5 and g(−3) means you have to put x= -3, so, g(−3)= 4(-3) -5 g(−3)= -12-5 g(−3)= -17